# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.2.1 The approximate problem
For any integer $k\geq 1$, let $h_{k}$ be a globally Lipschitz continuous function satisfying
@@
For $\widetilde{\mu}_{0}=-\Delta\varphi_{0}+\varPsi_{0}^{\prime}(\varphi_{0})$, we take the cut-off $\widetilde{\mu}_{0,k}=h_{k}\circ\widetilde{\mu}_{0}$. Since $\mu_{0}\in H^{1}(\Omega)$ and $\varphi_{0}\in H^{2}_{N}(\Omega)$, it follows from the superposition principle that
@@
Consider the nonlinear elliptic problem
$-\Delta\varphi_{0,k}+\varPsi_{0}^{\prime}(\varphi_{0,k})=\widetilde{\mu}_{0,k}\quad\text{in}\ \Omega,\qquad\partial_{\bm{n}}\varphi_{0,k}=0\quad\text{on}\ \partial\Omega.$
Thanks to [23, Lemma A.1], it admits a unique solution $\varphi_{0,k}\in H^{2}_{N}(\Omega)$ with $\varPsi_{0}^{\prime}(\varphi_{0,k})\in L^{2}(\Omega)$. Furthermore, we have
@@
Besides, we can infer from [11, Lemma A.1] that $\|\varPsi_{0}^{\prime}(\varphi_{0,k})\|_{L^{\infty}}\leq\|\widetilde{\mu}_{0,k}\|_{L^{\infty}}\leq k$, which together with the assumption (H2) implies that $\|\varphi_{0,k}\|_{C(\overline{\Omega})}\leq 1-\delta_{k}$ for some constant $\delta_{k}\in(0,1)$ depending on $k$. From the standard elliptic estimate, we also find $\varphi_{0,k}\in H^{3}(\Omega)$.
Next, for any positive integer $k$, we denote by $\sigma_{0,k}$ the unique solution of the linear elliptic problem
@@
$-\frac{1}{k}\Delta\sigma_{0,k}+\sigma_{0,k}=\sigma_{0}\qquad\text{in}\ \Omega,\qquad\partial_{\bm{n}}\sigma_{0,k}=0\quad\text{on}\ \partial\Omega.$
Since $\sigma_{0}\in H^{1}(\Omega)$, then it follows that
@@
$\sigma_{0,k}\in H^{2}_{N}(\Omega)\cap H^{3}(\Omega),\quad\overline{\sigma_{0,k}}=\overline{\sigma_{0}}\quad\text{and}\quad\|\sigma_{0,k}\|_{H^{1}}\leq\|\sigma_{0}\|_{H^{1}}.$
Moreover, we have $\|\sigma_{0,k}-\sigma_{0}\|\to 0$ as $k\to+\infty$.
Finally, the sequence of approximating velocity field $\{\bm{v}^{k}\}\subset C^{\infty}_{0}((0,+\infty);D(\bm{S}))$ is given as in the previous section.
@@
For any given $\gamma\in(0,1)$ and integer $k\geq\widehat{k}$ (thus the initial data are uniformly bounded with respective to $k$), we consider the following regularized problem
@@
$\partial_{t}\varphi^{\gamma,k}+\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}=\Delta\mu^{\gamma,k}-\alpha(\overline{\varphi^{\gamma,k}}-c_{0}),$ (3.36a)
$\mu^{\gamma,k}=-\Delta\varphi^{\gamma,k}+\varPsi^{\prime}(\varphi^{\gamma,k})-\chi\sigma^{\gamma,k}+\beta\mathcal{N}(\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}})+\gamma\partial_{t}\varphi^{\gamma,k},$ (3.36b)
$\partial_{t}\sigma^{\gamma,k}+\bm{v}^{k}\cdot\nabla\sigma^{\gamma,k}=\Delta(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k}),$ (3.36c)
@@
${\partial}_{\bm{n}}\varphi^{\gamma,k}={\partial}_{\bm{n}}\mu^{\gamma,k}={\partial}_{\bm{n}}\sigma^{\gamma,k}=0,\qquad\qquad$ on $\ \partial\Omega\times(0,+\infty),$ (3.37)
$\varphi^{\gamma,k}|_{t=0}=\varphi_{0,k}(x),\ \ \sigma^{\gamma,k}|_{t=0}=\sigma_{0,k}(x),\qquad$ in $\ \Omega,$ (3.38)
where $\varphi_{0,k}$, $\sigma_{0,k}$ are functions constructed above.
@@
By extending the results [21, Theorem A.1] (for the single viscous convective Cahn-Hilliard equation) and [28, Theorem 2.2] (for the coupled viscous Cahn–Hilliard-diffusion system without convection), we can establish the existence and uniqueness of a global strong solution $(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})$ to problem (3.36)–(3.38) on $[0,T]$ for any given $T>0$, that is,
@@
Assume that $\gamma\in(0,1)$ and $k\in\mathbb{Z}^{+}$ sufficiently large. Let $T>0$. The approximate problem (3.36)–(3.38) admits a unique strong solution $(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})$ on $[0,T]$ such that
@@
$\mu^{\gamma,k}\in L^{\infty}(0,T;H^{2}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega)),$
$\sigma^{\gamma,k}\in L^{\infty}(0,T;H^{2}(\Omega))\cap L^{2}(0,T;H^{3}(\Omega)),$
$\partial_{t}\sigma^{\gamma,k}\in L^{\infty}(0,T;L^{2}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega)).$
The triple $(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})$ satisfies the system (3.36) almost everywhere in $\Omega\times(0,T)$, the boundary condition (3.37) almost everywhere on $\partial\Omega\times(0,T)$, and the initial condition (3.38) in $\Omega$. Moreover, there exists some $\widehat{\delta}_{k}\in(0,\delta_{k})$ depending on $\gamma$ such that
$\|\varphi^{\gamma,k}(t)\|_{C(\overline{\Omega})}\leq 1-\widehat{\delta}_{k},\quad\forall\,t\in[0,T].$
3.2.2 Uniform estimates
In what follows, we derive estimates for $(\varphi^{\gamma,k},\mu^{\gamma,k},\sigma^{\gamma,k})$ that are independent of the parameters $\gamma$, $k$ and even $T$.
First estimate. Similar to (3.14) and (3.15), we have
@@ 3.2.2 Uniform estimates
and
$\overline{\sigma^{\gamma,k}}(t)=\overline{\sigma^{\gamma,k}}(0)=\overline{\sigma_{0,k}},\quad\forall\,t\in[0,T].$ (3.40)
Second estimate. Similar to (3.16), we find
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}^{\gamma,k}(t)+\widetilde{\mathcal{D}}^{\gamma,k}(t)$ $=-\alpha\big{(}\overline{\varphi^{\gamma,k}}-c_{0}\big{)}\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x-\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\mu^{\gamma,k}\,\mathrm{d}x$
$\quad+\chi\int_{\Omega}(\bm{v}^{k}\cdot\nabla\sigma^{\gamma,k})\varphi^{\gamma,k}\,\mathrm{d}x,$ (3.41)
for almost all $t\in(0,T)$, where
@@
$\mathcal{F}^{\gamma,k}(t)$ $=\int_{\Omega}\Big{(}\frac{1}{2}|\nabla\varphi^{\gamma,k}(t)|^{2}+\varPsi(\varphi^{\gamma,k}(t))-\chi\sigma^{\gamma,k}(t)\varphi^{\gamma,k}(t)+\frac{1}{2}|\sigma^{\gamma,k}(t)|^{2}\Big{)}\,\mathrm{d}x$
$\quad+\frac{\beta}{2}\|\nabla\mathcal{N}\big{(}\varphi^{\gamma,k}(t)-\overline{\varphi^{\gamma,k}(t)}\big{)}\|^{2},$
$\widetilde{\mathcal{D}}^{\gamma,k}(t)$ $=\|\nabla\mu^{\gamma,k}(t)\|^{2}+\|\nabla(\sigma^{\gamma,k}(t)-\chi\varphi^{\gamma,k}(t))\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}.$
As a consequence, it holds
$\|\varphi^{\gamma,k}(t)\|_{H^{1}}^{2}+\|\sigma^{\gamma,k}(t)\|^{2}+\int_{0}^{t}\big{(}\|\nabla\mu^{\gamma,k}(s)\|^{2}+\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})(s)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\big{)}\,\mathrm{d}s$
where $C>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, $\max_{r\in[-1,1]}|\varPsi(r)|$, $\|\varphi_{0}\|_{H^{1}}$, $\|\sigma_{0}\|$, coefficients of the system and $\Omega$, but it is independent of $T$, $\gamma$ and $k$.
Third estimate. We proceed to derive $L^{q}$-estimate for $\sigma^{\gamma,k}$. Testing (3.36c) by $(\sigma^{\gamma,k})^{3}$ yields
@@
$\frac{1}{4}\frac{\mathrm{d}}{\mathrm{d}t}\|\sigma^{\gamma,k}\|_{L^{4}}^{4}+\frac{3}{4}\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}=-\int_{\Omega}\chi\Delta\varphi^{\gamma,k}(\sigma^{\gamma,k})^{3}\,\mathrm{d}x.$ (3.43)
From the Sobolev embedding theorem and Poincaré’s inequality, it follows that
$\|\sigma^{\gamma,k}\|_{L^{12}}^{2}=\|(\sigma^{\gamma,k})^{2}\|_{L^{6}}\leq C\left(\|\nabla(\sigma^{\gamma,k})^{2}\|+\int_{\Omega}(\sigma^{\gamma,k})^{2}\,\mathrm{d}x\right).$ (3.44)
By (3.42), (3.44) and Young’s inequality, we get
$\|(\sigma^{\gamma,k})^{3}\|_{L^{\frac{6}{5}}}^{2}=\|\sigma^{\gamma,k}\|_{L^{\frac{18}{5}}}^{6}$ $\leq\|\sigma^{\gamma,k}\|^{\frac{14}{5}}\|\sigma^{\gamma,k}\|_{L^{12}}^{\frac{16}{5}}\leq\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}+C.$
The above estimate combined with Hölder’s and Young’s inequalities yields
@@
$-\int_{\Omega}\chi\Delta\varphi^{\gamma,k}(\sigma^{\gamma,k})^{3}\,\mathrm{d}x$ $\leq C|\chi|\|\Delta\varphi^{\gamma,k}\|_{L^{6}}\|(\sigma^{\gamma,k})^{3}\|_{L^{\frac{6}{5}}}$
$\leq C\chi^{2}\|\Delta\varphi^{\gamma,k}\|^{2}_{L^{6}}+\frac{1}{4}\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}+C.$ (3.45)
Hence, it follows from (3.43) and (3.45) that
@@
$\frac{1}{4}\frac{\mathrm{d}}{\mathrm{d}t}\|\sigma^{\gamma,k}\|_{L^{4}}^{4}+\frac{1}{2}\|\nabla(\sigma^{\gamma,k})^{2}\|^{2}\leq C\chi^{2}\|\Delta\varphi^{\gamma,k}\|^{2}_{L^{6}}+C.$ (3.46)
@@
To handle the first term on the right-hand side, we consider the following elliptic problem with singular nonlinearity:
@@
$\begin{cases}-\Delta\varphi^{\gamma,k}+\varPsi_{0}^{\prime}(\varphi^{\gamma,k})=\mu^{\gamma,k}+\theta_{0}\varphi^{\gamma,k}+\chi\sigma^{\gamma,k}-\gamma\partial_{t}\varphi^{\gamma,k},\quad&\text{ in }\Omega,\\ \partial_{\bm{n}}\varphi^{\gamma,k}=0,\quad&\text{ on }\partial\Omega.\end{cases}$
According to the argument in [22, Section 4] (see also [1, 11]), it holds

### v2

@@ 3.2.2 Decoupled auxiliary problem for $(\varphi,\mu)$
The regularized problem. We first approximate the initial data. For any integer $k\geq 1$, let $h_{k}$ be a globally Lipschitz continuous function satisfying
@@
Define
$\widetilde{\mu}_{0}=-\Delta\varphi_{0}+\varPsi_{0}^{\prime}(\varphi_{0})=\mu_{0}+\theta_{0}\varphi_{0}.$
We take the cut-off $\widetilde{\mu}_{0,k}=h_{k}\circ\widetilde{\mu}_{0}$. Since $\mu_{0}\in H^{1}(\Omega)$ and $\varphi_{0}\in H^{2}_{N}(\Omega)$, it follows from the superposition principle that
@@
Given $\widetilde{\mu}_{0,k}\in H^{1}(\Omega)$, let us consider the nonlinear elliptic problem
$\begin{cases}-\Delta\varphi_{0,k}+\varPsi_{0}^{\prime}(\varphi_{0,k})=\widetilde{\mu}_{0,k},\quad\text{in}\ \Omega,\\ \partial_{\bm{n}}\varphi_{0,k}=0,\qquad\qquad\qquad\quad\ \ \text{on}\ \partial\Omega.\end{cases}$
Thanks to [28, Lemma A.1], it admits a unique solution $\varphi_{0,k}\in H^{2}_{N}(\Omega)$ with $\varPsi_{0}^{\prime}(\varphi_{0,k})\in L^{2}(\Omega)$. Furthermore, we have
@@
Besides, we infer from [14, Lemma A.1] that $\|\varPsi_{0}^{\prime}(\varphi_{0,k})\|_{L^{\infty}}\leq\|\widetilde{\mu}_{0,k}\|_{L^{\infty}}\leq k$, which together with the assumption (H2) implies the separation property
$\|\varphi_{0,k}\|_{C(\overline{\Omega})}\leq 1-\delta_{k},$
for some constant $\delta_{k}\in(0,1)$ depending on $k$. This yields $\varPsi_{0}^{\prime}(\varphi_{0,k})\in H^{1}(\Omega)$. Using the standard elliptic estimate, we further obtain $\varphi_{0,k}\in H^{3}(\Omega)$.
Next, we approximate the given velocity field $\bm{v}$ and the source term $\sigma$. Like in Section 3.1.1, we take a convergent sequence $\big\{\bm{v}^{k}\big\}\subset C^{\infty}_{0}((0,+\infty);D(\bm{S}))$ satisfying (3.8). Concerning $\sigma$, for any $k\in\mathbb{Z}^{+}$, we denote by $\sigma^{k}$ the (unique) solution of the following linear elliptic problem
@@
$\begin{cases}\dfrac{1}{k}(-\Delta\sigma^{k}+\sigma^{k})+\sigma^{k}=\sigma,\quad\ \ \text{in}\ \Omega,\\ \partial_{\bm{n}}\sigma^{k}=0,\qquad\qquad\qquad\qquad\text{on}\ \partial\Omega.\end{cases}$
The operator $\Delta+I$ subject to the homogeneous Neumann boundary condition is an canonical isomorphism from $H^{1}(\Omega)\to(H^{1}(\Omega))^{\prime}$ due to the Riesz theorem. In view of the assumption on $\sigma$, we infer from the elliptic regularity theory that
$\sigma^{k}\in L^{2}_{\mathrm{uloc}}([0,+\infty);H^{3}(\Omega)\cap H^{2}_{N}(\Omega))\cap L^{\infty}(0,+\infty;W^{2,6}(\Omega))\cap H^{1}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega)).$
By interpolation, we also find $\sigma^{k}\in BUC([0,+\infty);H^{2}(\Omega))$, while from the assumption $\sigma(0)\in H^{1}(\Omega)$, we obtain $\sigma^{k}(0)\in H^{3}(\Omega)\cap H^{2}_{N}(\Omega)$. Applying Lemma A.2 with $V=H^{1}(\Omega)$, $H=L^{2}(\Omega)$, $\varepsilon^{2}=1/k$, $J=-\Delta+I$ for $u=\sigma$ and then $u=\partial_{t}\sigma$, we can conclude that
$\|\sigma^{k}\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega))}\leq\|\sigma\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega))},$
$\|\partial_{t}\sigma^{k}\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime})}\leq\|\partial_{t}\sigma\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime})},$
$\|\sigma^{k}(0)\|_{H^{1}}\leq\|\sigma(0)\|_{H^{1}},$
moreover, as $k\to+\infty$, it holds
$\sigma^{k}\to\sigma\quad\text{in}\ \ L^{2}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega)),$
@@
$\partial_{t}\sigma^{k}\to\partial_{t}\sigma\quad\text{in}\ \ L^{2}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime}),$
$\sigma^{k}(0)\to\sigma(0)\quad\text{in}\ \ H^{1}(\Omega).$
Testing the equation of $\sigma^{k}$ by $(\sigma^{k})^{q}$, $q=1,3,5$, integrating over $\Omega$, using Hölder’s inequality and Young’s inequality, we obtain
$\underbrace{\frac{1}{k}\int_{\Omega}\left(q(\sigma^{k})^{q-1}|\nabla\sigma^{k}|^{2}+(\sigma^{k})^{q+1}\right)\,\mathrm{d}x}_{\geq 0}+\|\sigma^{k}\|_{L^{q+1}}^{q+1}$ $=\int_{\Omega}\sigma(\sigma^{k})^{q}\,\mathrm{d}x$
$\leq\|\sigma\|_{L^{q+1}}\|\sigma^{k}\|_{L^{q+1}}^{q}$
$\leq\frac{1}{q+1}\|\sigma\|_{L^{q+1}}^{q+1}+\frac{q}{q+1}\|\sigma^{k}\|_{L^{q+1}}^{q+1},$
which implies the uniform bounds
$\|\sigma^{k}\|_{L^{\infty}(0,+\infty;L^{q+1}(\Omega))}\leq\|\sigma\|_{L^{\infty}(0,+\infty;L^{q+1}(\Omega))},\quad q=1,3,5.$
@@
For any given $\gamma\in(0,1)$ and integer $k\geq\widehat{k}$ (thus the approximated initial data are uniformly bounded with respect to $k$), we consider the following regularized problem
@@
$\partial_{t}\varphi^{\gamma,k}+\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}=\Delta\mu^{\gamma,k}-\alpha(\overline{\varphi^{\gamma,k}}-c_{0}),$ in $\ \Omega\times(0,+\infty),$ (3.66a)
$\mu^{\gamma,k}=-\Delta\varphi^{\gamma,k}+\varPsi^{\prime}(\varphi^{\gamma,k})-\chi\sigma^{k}$
@@
$\qquad\quad+\beta\mathcal{N}(\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}})+\gamma\partial_{t}\varphi^{\gamma,k},$ in $\ \Omega\times(0,+\infty),$ (3.66b)
${\partial}_{\bm{n}}\varphi^{\gamma,k}={\partial}_{\bm{n}}\mu^{\gamma,k}=0,\qquad\qquad$ on $\ \partial\Omega\times(0,+\infty),$ (3.66c)
$\varphi^{\gamma,k}|_{t=0}=\varphi_{0,k}(x),\qquad$ in $\ \Omega.$ (3.66d)
@@
By extending the previous results [25, Theorem A.1] (for the single convective viscous Cahn–Hilliard equation) and [32, Theorem 2.2] (for the coupled viscous Cahn–Hilliard–diffusion system without convection), we can establish the existence and uniqueness of a global strong solution $(\varphi^{\gamma,k},\mu^{\gamma,k})$ to problem (3.66) on $[0,T]$ for any given $T\geq 1$, that is,
@@
Assume that $\Omega$ is a bounded domain in $\mathbb{R}^{3}$ with boundary $\partial\Omega$ of class $C^{3}$, (H2)–(H3) are satisfied, $\gamma\in(0,1)$ and $k\in\mathbb{Z}^{+}$ is sufficiently large. For any $T\geq 1$, the regularized problem (3.66) admits a unique strong solution $(\varphi^{\gamma,k},\mu^{\gamma,k})$ on $[0,T]$ such that
@@
$\mu^{\gamma,k}\in L^{\infty}(0,T;H^{2}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega)).$
The pair $(\varphi^{\gamma,k},\mu^{\gamma,k})$ satisfies the equation (3.66a)–(3.66b) almost everywhere in $\Omega\times(0,T)$, the boundary condition (3.66c) almost everywhere on $\partial\Omega\times(0,T)$, and the initial condition (3.66d) in $\Omega$. Moreover, there exists some constant $\delta_{\gamma,k}\in(0,\delta_{k})$ depending on $\gamma$ such that
$\|\varphi^{\gamma,k}(t)\|_{C(\overline{\Omega})}\leq 1-\delta_{\gamma,k},\quad\forall\,t\in[0,T].$ (3.67)
After the above preparations, we are now in a position to prove Proposition 3.2-(2).
Proof of Proposition 3.2-(2).
The uniqueness of strong solutions is a direct consequence of the uniqueness result for weak solutions established in Proposition 3.2-(1). To establish the existence, we derive estimates for $(\varphi^{\gamma,k},\mu^{\gamma,k})$ that are independent of the approximating parameters $\gamma$, $k$ and time.
First estimate. Analogously to (3.60), we have the mass relation
@@
Second estimate. A similar argument for (3.61) yields the following equality
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}\|\nabla\varphi^{\gamma,k}\|^{2}+\int_{\Omega}\varPsi(\varphi^{\gamma,k})\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}})\|^{2}\right)+\|\nabla\mu^{\gamma,k}\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|^{2}$
@@
$\quad=-\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})(\mu^{\gamma,k}+\chi\sigma^{k})\,\mathrm{d}x-\chi\int_{\Omega}\nabla\mu^{\gamma,k}\cdot\nabla\sigma^{k}\,\mathrm{d}x$
@@
$\qquad-\alpha(\overline{\varphi^{\gamma,k}}-c_{0})\int_{\Omega}(\mu^{\gamma,k}+\chi\sigma^{k})\,\mathrm{d}x,$ (3.69)
for almost all $t\in(0,T)$. We can estimate the right-hand side of (3.69) like before, keeping in mind possible modifications due to the additional viscous term $\gamma\partial_{t}\varphi^{\gamma,k}$ in $\mu^{\gamma,k}$. The first two terms can be handled as in (3.62). For the last term, recalling the estimates [34, (3.15), (3.16)] and using the bound $\|\varphi^{\gamma,k}\|_{L^{\infty}}\leq 1$ (cf. (3.67)), we infer that
@@
$\left|\int_{\Omega}\mu^{\gamma,k}\,\mathrm{d}x\right|$ $\leq C\big(\|\nabla\mu^{\gamma,k}\|+\|\sigma^{k}-\overline{\sigma^{k}}\|+\|\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big)\|\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}}\|+C$
$\leq C\big(\|\nabla\mu^{\gamma,k}\|+\|\nabla\sigma^{k}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|+1\big).$ (3.70)
Then from (3.69) we can deduce that
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}\|\nabla\varphi^{\gamma,k}\|^{2}+\int_{\Omega}\varPsi(\varphi^{\gamma,k})\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}})\|^{2}+C_{7}\right)+\frac{1}{4}\|\nabla\mu^{\gamma,k}\|^{2}+\frac{\gamma}{4}\|\partial_{t}\varphi^{\gamma,k}\|^{2}$
$\qquad+C_{10}\left(\frac{1}{2}\|\nabla\varphi^{\gamma,k}\|^{2}+\int_{\Omega}\varPsi(\varphi^{\gamma,k})\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}})\|^{2}+C_{7}\right)$
$\quad\leq C_{11}(\|\bm{v}^{k}\|^{2}+\|\sigma^{k}\|_{H^{1}}^{2}+1),$ (3.71)
for some positive constants $C_{10},C_{11}$ that are independent of $\gamma$, $k$ and time. Here, the constant $C_{7}>0$ is the same as in (3.63). Based on the construction of approximate data $\bm{v}^{k}$, $\sigma^{k}$, and $\varphi_{0,k}$, the differential inequality (3.71) combined with Lemma A.1 yields the following estimate:
$\sup_{t\in[0,T]}\left(\frac{1}{2}\|\nabla\varphi^{\gamma,k}(t)\|^{2}+\int_{\Omega}\varPsi(\varphi^{\gamma,k}(t))\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{\gamma,k}-\overline{\varphi^{\gamma,k}})(t)\|^{2}\right)$
$\quad+\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\nabla\mu^{\gamma,k}(s)\|^{2}\,\mathrm{d}s+\sup_{t\in[0,T-1]}\int_{t}^{t+1}\gamma\|\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\,\mathrm{d}s\leq C,$ (3.72)
where the positive constant $C$ depends on $\Omega$, the coefficients of the system, $\max_{r\in[-1,1]}|\varPsi(r)|$, $\overline{\varphi_{0}}$, $\|\varphi_{0}\|_{H^{1}}$, $\|\bm{v}\|_{L^{2}(0,+\infty,\bm{L}^{2}(\Omega))}$, $\sup_{t\geq 0}\int_{t}^{t+1}\|\sigma(s)\|_{H^{1}(\Omega)}^{2}\,\mathrm{d}s$, but it is independent of $\gamma$, $k$ and $T$.
@@
Third estimate. Consider the following elliptic problem with singular nonlinearity:
@@
$\begin{cases}-\Delta\varphi^{\gamma,k}+\varPsi_{0}^{\prime}(\varphi^{\gamma,k})=\mu^{\gamma,k}+\theta_{0}\varphi^{\gamma,k}+\chi\sigma^{k}-\gamma\partial_{t}\varphi^{\gamma,k},\quad&\text{ in }\Omega,\\ \partial_{\bm{n}}\varphi^{\gamma,k}=0,\quad&\text{ on }\partial\Omega.\end{cases}$
According to the argument in [26, Section 4] (see also [1, 14]), it holds